Futuristic was a newly setup firm in 2019. It is in the business of providing artwork. However, being in the business of standard art objects (non-essential items), it did not do well once it was hit by COVID-19. It has a selling price of ₹2,500 per piece and a variable cost of ₹1,000 per piece. It incurs an annual fixed cost of ₹30,00,000. a. The manager of the business wishes to know-  the number of art pieces they must sell to be at the break-even point  the contribution margin whether the firm is earning a profit or incurring a loss by selling 2,100 art pieces



1
Expert's answer
2021-11-07T14:01:21-0500

a. The break-even point is:

"BEP = \frac{3,000,000} {2,500 - 1,000} = 2,000"

units.

The contribution margin is:

"CM = \frac{2,500-1,000} {2,500} = 0.6" or 60%.

b. If Q = 2,100, then the total profit is:

"TP = 2,500\u00d72,100 - (3,000,000 + 1,000\u00d72,100) = 150,000."